OpenStudy (ashy98):

Can some one help me ??(:

OpenStudy (texaschic101):

question please

OpenStudy (hwyl):

YES

OpenStudy (hwyl):

THE ANSWER IS YES

OpenStudy (anonymous):

XD

OpenStudy (ashy98):

What is the simplified form of \[\sqrt[7]{x}*\sqrt[7]{x}*\sqrt[7]{x}*\sqrt[7]{x}\]

OpenStudy (ashy98):

I got \[\sqrt[7]{x4}\]

OpenStudy (michele_laino):

hint: \[\huge \sqrt[7]{x} = {x^{\frac{1}{7}}}\]

OpenStudy (michele_laino):

correct!

OpenStudy (ashy98):

Thank you !!(:

OpenStudy (michele_laino):

:)